40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Dec 7 13:11:57 2014, manasa, Summary, General, Frequency Offset Locking - To Do List FOLtodolist.pdf
    Reply  Mon Dec 8 15:54:39 2014, manasa, Summary, General, Dec 8 - Check Frequency Counter module 
       Reply  Mon Dec 8 20:53:51 2014, diego, Summary, General, Dec 8 - Check Frequency Counter module 
          Reply  Tue Dec 9 16:06:46 2014, diego, Summary, General, Dec 8 - Check Frequency Counter module 
    Reply  Tue Dec 9 00:30:27 2014, manasa, Summary, General, Dec 9 - Elaborate to do list FOL_FiberChassis.pdf
       Reply  Tue Dec 9 16:07:16 2014, manasa, Summary, General, Dec 9 - FC module and fiber chassis FCmodule.png
    Reply  Wed Dec 10 16:12:29 2014, manasa, Summary, General, Dec 10 - PSL table FiberMod_front.jpgFOL_fiber.jpg
       Reply  Thu Dec 11 12:50:12 2014, manasa, Summary, General, PSL table optical layout 
          Reply  Thu Dec 11 23:34:06 2014, manasa, Summary, General, PSL table optical layout 
             Reply  Fri Dec 12 02:30:25 2014, Jenne, Summary, General, PSL table optical layout 
    Reply  Fri Dec 12 14:38:39 2014, manasa, Summary, General, Frequency Offset Locking - To Do List (Revised) FOLtodolist.pdf
       Reply  Fri Dec 12 15:19:04 2014, manasa, Summary, General, Dec 12 - PSL table PSLtoFiber.pngPSLfiberChassis.pngPSLfiberChassis.png
       Reply  Wed Dec 17 11:57:56 2014, manasa, Summary, General, Y arm optical layout 58.png
Message ID: 10775     Entry time: Wed Dec 10 16:12:29 2014     In reply to: 10760     Reply to this: 10780
Author: manasa 
Type: Summary 
Category: General 
Subject: Dec 10 - PSL table 

Quote:

Attached is the timeline for Frequency Offset Locking related activities. All activities will be done mostly in morning and early afternoon hours.

I was working around the PSL table today.

I wanted to modify the telescope that couples PSL light into the fiber; now that I have the translation stages for the lenses. I could not finish it as the locking work started earlier than usual this afternoon. I measured the out of loop noise for ALS error signals before I opened the PSL enclosure. X and Y beat notes were at -18dBm at 49.3MHz and -29.56dBm at 62.2MHz for this measurement. DTT data can be found in /users/manasa/data/141210/ALSoutLoop.xml; so there is reference to go back to in case of any damage done due to the work on the PSL table.

Also, I received the front and back panels for the Fiber chassis and put it together. Find photos (front panel and inside) of chassis in attachment. This will go inside the PSL enclosure tomorrow.

FiberMod_front.jpg    FOL_fiber.jpg

 

ELOG V3.1.3-